# Bayesian Adaptive Tucker Decompositions for Tensor Factorization

### v1

@@ Abstract
Tucker tensor decomposition offers a more effective representation for multiway data compared to the widely-used PARAFAC model. However, its flexibility brings the challenge of selecting the appropriate latent multi-rank. To overcome the issue of pre-selecting the latent multi-rank, we introduce a Bayesian adaptive Tucker decomposition model that infers the multi-rank automatically via an infinite increasing shrinkage prior. The model introduces local sparsity in the core tensor, inducing rich and at the same time parsimonious dependency structures. Posterior inference proceeds via an efficient adaptive Gibbs sampler, supporting both continuous and binary data and allowing for straightforward missing data imputation when dealing with incomplete multiway data. We discuss fundamental properties of the proposed modeling framework, providing theoretical justification. Simulation studies and applications to chemometrics and complex ecological data offer compelling evidence of its advantages over existing tensor factorization methods.
@@ 1 Introduction
In many applications, data naturally come in the form of multiway arrays or tensors such as recommender system (Bi et al., 2017), ecology (Frelat et al., 2017), genomics (Hore et al., 2016) and chemometrics (Tomasi and Bro, 2005). In a general setting, let $\mathcal{Y}=\{y_{i_{1}\cdots i_{K}},\ i_{k}=1,\dots,n_{k}\}$ denote a $K$-way tensor and we assume that $\mathcal{Y}\in\mathbb{R}^{n_{1}\times\dots\times n_{k}}$, with the integer $n_{k}$ being the dimension of $\mathcal{Y}$ along the $k$-th way. We are motivated by settings where $K$ is relatively small. For instance, in ecological studies examining species co-occurrence, $K=3$ with data organized in a three-way array structured as site $\times$ species $\times$ year. One naive approach to analyze this type of data is to flatten the data into a matrix and then use existing methods for matrix factorization. However, this transformation fails to preserve the data structure, making learning low-dimensional relationships more difficult.
@@
Tensor decomposition methods based on the low-rank tensor approximation (Kolda and Bader, 2009) offer a powerful tool to analyze multiway data, by providing a concise representation that effectively captures the dominant underlying features. The two main forms of tensor decomposition, which extend the matrix singular value decomposition, are the Tucker decomposition and its restriction known as the parallel factor (PARAFAC) decomposition. PARAFAC factorization decomposes a tensor into a sum of $R$ rank-one tensor, where $R$ is referred to as the rank of the tensor, while Tucker factorization decomposes a tensor into a small core tensor multiplied by a matrix along each mode. Tucker representation is more flexible than the PARAFAC, as it includes a multi-rank $(R_{1},\dots,R_{k})$ that allows for different decomposition ranks for various tensor dimension margins. This flexibility enables a more effective data compression, as not all the dimensions of a tensor might require all $R$ principal components in order to be faithfully represented by such a decomposition (Wang et al., 2005; Kim and Choi, 2007). In this case, PARAFAC rank can be considerably higher than the $K$ ranks composing the Tucker multi-rank. Tucker representation is particularly effective when the tensor dimensions $n_{1},\dots,n_{K}$ differ from each other.
@@
Bayesian tensor decomposition models (Chu and Ghahramani, 2009) are naturally appealing, as they allow for straightforward quantification of uncertainty in estimating model parameters, while providing a principled mechanism for dealing with missing data. Partially observed multiway data, indeed, are abundant in numerous fields; see for example Tomasi and Bro (2005) for applications in chemometrics. Missing data can arise in a variety of settings due to loss of information, errors in the data collection process, or costly experiments. The problem of recovering latent structures from incomplete tensors to reconstruct the missing entries is often referred to as tensor completion. For tensor factorization models in a Bayesian framework see for instance Hoff (2011) and Xu et al. (2013). Refer to Yiyao and Weining (2024) for a general review of Bayesian methods in tensor analysis, including topics beyond tensor completion, such as tensor regression models (Guhaniyogi et al., 2017; Papadogeorgou et al., 2021).
@@
An open problem in low-rank tensor factorization literature is the choice of the latent rank. Current practice typically specifies the ranks in advance or uses heuristics for choosing the ranks based on repeated fitting under different choices. However, in the Tucker case, an exhaustive search over the space of possible multi-ranks may be infeasible, since the number of possibilities increases exponentially with the order of the tensor. A more elegant approach can be found in Bayesian nonparametric methods with Rai et al. (2014) and Zhao et al. (2015) being interesting contributions in the PARAFAC context.
@@
Motivated by the above considerations and the need for an intermediate formulation that balances the simplicity of the PARAFAC with the flexibility of the Tucker in a parsimonious way, we introduce a flexible Bayesian Tucker decomposition model where the multi-rank is not pre-specified but is adaptively learned from the data. We rely on successful Bayesian nonparametric models that include more than enough latent elements, with their selection induced by appropriate shrinkage priors that adaptively eliminate unnecessary components. Exploiting the literature on infinite factor model (Bhattacharya and Dunson, 2011; Legramanti et al., 2020; Schiavon et al., 2022) we introduce a multiway extension of the cumulative shrinkage process (CUSP) of Legramanti et al. (2020) to apply joint shrinkage for multiple latent quantities along different dimensions. Besides providing shrinkage towards low-rank decomposition of the tensor margins, the proposed method also enables local shrinkage in the latent factors by assuming a suitable shrinkage prior for the elements of the core tensor. Incorporating zero entries in the core tensor is crucial for reducing complexity, as inference in the Tucker model scales with the size of the core tensor. By shrinking towards zero the elements of the core tensor we are moving it to something similar to a superdiagonal tensor with different dimension margins, and thus bridging PARAFAC and Tucker decompositions. For a contribution in this direction for probability tensors for multivariate categorical data see the collapsed Tucker decomposition in Johndrow et al. (2017).
@@
In the next section, we will first establish the notation, introduce the basic formulation of our Bayesian adaptive Tucker (AdaTuck) decomposition and discuss appealing theoretical properties. We will outline straightforward implementations for modeling both continuous and binary incomplete multiway data using fully conjugate Bayesian analysis. Posterior distributions are approximated through closed-form Gibbs sampling updates, making inference highly scalable and straightforward to predict missing entries in the data, as discussed in Section 3. The empirical performances of the proposed approach are tested against alternatives in Section 4, including simulated and real data examples. Specifically, we will analyze data from a chemometrics application and a complex ecological study.
@@ 2.2 Bayesian model specification
Proposition 1 provides a simple and intuitive formulation for the expected value of the latent multi-rank, yielding useful insights for the choice of the hyperparameters $\alpha^{(k)}$.
@@
Let $M_{\textbf{i}}^{R_{1},\dots,R_{k}}=\sum_{\mathbf{r},r_{k}=R_{k}+1}^{\infty}g_{\mathbf{r}}\prod_{k=1}^{K}u^{(k)}_{i_{k}r_{k}}$ be the residuals for a generic index vector $\mathbf{i}$. Assume $E(g_{\mathbf{r}}^{2})\leq\zeta$ with $\zeta$ finite. Then, $\forall\epsilon>0$,
$\mathrm{pr}\{(M_{\textbf{i}}^{R1,\dots,R_{k}})^{2}>\epsilon\}<\zeta\prod_{k=1}^{K}\frac{1}{\epsilon\alpha^{(k)}\{\alpha^{(k)}(1+\alpha^{(k)})^{-1}\}^{R_{k}}}.$
@@
The hierarchical Bayesian model specification is completed specifying a prior distribution for the error variance in (3). For simplicity we let $\sigma^{2}\sim\text{InvGa}(a_{\sigma},b_{\sigma})$. When $\mathcal{Y}$ has $m<n$ missing entries, the likelihood can be easily adapted from Equation (3) by removing, from the covariance matrix, the associated $m$ rows and columns. In this case, our Bayesian approach seamlessly allows to perform missing data imputation by simulating from the posterior predictive distribution. See Sections 3 and 4.1 for details.
@@ 2.4 Consistency in rank selection
One of the key aspects of AdaTuck is its ability to learn the rank along each dimension, through the multiway increasing shrinkage prior. Consistent with this characteristic, in this section we investigate the asymptotic properties of the adopted multiway CUSP in a simplified setting. This analysis extend the discussion in Section 3.3 of Kowal and Canale (2023) and relates to the problem of estimating a high-dimensional mean vector from a single multivariate observation (Castillo and van der Vaart, 2012; Ročková, 2018).
@@
with $\theta^{(k)}_{i}$ following the prior structure in (4). While this setting differs from the Tucker decomposition model, we observe that the above likelihood shares some similarities with the proposed models as it corresponds to the multiplicative term in the definition of the elementwise Tucker decomposition in (2). Hence $y_{i}\sim N(\eta_{i},1)$ for $i=1,2,\dots$ with $\eta_{i}=\prod_{k=1}^{K}u_{i}^{(k)}$. We assume that the true data generating processing is
@@
According to Lemma 1, the prior probability that $M_{n}$ exceeds a small threshold is exponentially small. To connect the prior behavior in Lemma 1 to the posterior we denote with $\mathrm{pr}_{0}(\cdot)$ and $E_{0}(\cdot)$ the probability and the expectation respectively under the true distribution of the data, described in (10). Theorem 2 shows that the posterior probability that $M_{n}$ exceeds a shrinking $\varepsilon_{n}$ for $n\to\infty$ goes to zero in expectation thus providing theoretical justification for the use of the proposed multiway CUSP to learn the ranks of model (1).
@@
Let $\epsilon_{n}\to 0$ with $\epsilon_{n}^{1/\tilde{R}^{0n}}>\alpha^{(k)}/(1+\alpha^{(k)})$ for any $k=1\dots,K$. We assume $A\tilde{R}^{0n}/n<1/2$ with $A>1/2$. Let $\theta_{\infty}=\theta_{\infty}(n)$ in the multiway CUSP prior (4) with $\theta_{\infty}(n)>r_{n}$ for $1<r<\sqrt{n}$. Then,
$\lim_{n\to\infty}E_{0}\{\mathrm{pr}(M_{n}>\varepsilon_{n}\mid y_{1},\dots,y_{n})\}=0.$
@@ 4.1 Synthetic data
    In this section, we evaluate the performance of AdaTuck with synthetic data. Specifically, we assess its performance in terms of learning the true latent multi-rank of the Tucker factorization, under correct specification. Additionally, we assess its goodness-of-fit by randomly removing some entries of the tensor data before fitting and then imputing these missing values by exploiting the posterior predictive distribution.
@@
    To assess the performance of our proposed approach in learning the true underlying multi-rank $(R_{1}^{0},R_{2}^{0},R_{3}^{0})$, we compute the Monte Carlo average over the $B$ replicates of the posterior median of $R_{1}^{*}$, $R_{2}^{*}$ and $R_{3}^{*}$ under the AdaTuck model. The results for the $B$ different datasets across the two different scenarios are shown in Figure 1. In both scenarios, our proposed method correctly recovers the true multi-rank for all tensor margins, thus efficiently learning the truncation level in the adaptive Gibbs sampler. AdaTuck posterior estimations of the multi-rank highly concentrate around the true values $(R_{1}^{0},R_{2}^{0},R_{3}^{0})$, and the results are robust with respect to the number of missing values in the data. Figure S1 in the Supplementary Materials illustrates the results for the probit-AdaTuck model in the same scenarios, showing its effectiveness in recovering the true multi-rank also in the binary data setting.
@@
    Subsequently, we evaluate the performance of the proposed method in predicting the missing values and reconstructing the original tensor for the two scenarios. We benchmark the results with the Bayesian tensor factorization method proposed in Khan and Ammad-ud din (2016) and implemented in the R package tensorBF, which targets multiway data with incomplete records. This method implements a Bayesian PARAFAC decomposition and achieves rank optimization through a sparsity parameter to remove redundant factors. We run tensorBF with the default options for the same number of iterations we used for AdaTuck. We compare the two methods by computing the Monte Carlo mean square error (MSE) in predicting the missing values with the reconstructed tensor based on the latent structure. Table 1 summarizes, for each scenario and model, the median and the interquartile range of the MSE obtained from the $B$ measures generated from different simulations. Across all scenarios AdaTuck outperforms the competitor in predicting the missing entries of the tensor. As expected, tensorBF, which implements a PARAFAC decomposition, performs better in the scenario similar to the PARAFAC structure. Instead, AdaTuck with its greater flexibility provides good results in both scenarios. Additionally, AdaTuck is robust to variations in the amount of available data, maintaining stable performance even as the percentage of missing values increases, whereas the competitor’s performance gets worse as missing entries increase.
    Table 1: Performance of AdaTuck and TensorBf in predicting the missing entries for the 20 simulated datasets for the two scenarios. The shorthand ‘NA(%)’ represents the percentage of missing entries with respect to the total tensor entries.
@@
    $(R_{1}^{0},R_{2}^{0},R_{3}^{0})$ NA (%) MSE (median) MSE (IQR) MSE (median) MSE (IQR)
    (5,5,5) 30 0.22 0.01 33.00 47.27
    50 0.23 0.02 46.73 374.78
    (10,7,3) 30 0.23 0.01 138.27 121.13
    50 0.24 0.01 146.28 153.58

### v2

@@ Abstract
Tucker tensor decomposition offers a more effective representation for multiway data compared to the widely used PARAFAC model. However, its flexibility brings the challenge of selecting the appropriate latent multi-rank. To overcome the issue of pre-selecting the latent multi-rank, we introduce a Bayesian adaptive Tucker decomposition model that infers the multi-rank automatically via an infinite increasing shrinkage prior. The model introduces local sparsity in the core tensor, inducing rich and at the same time parsimonious dependency structures. Posterior inference proceeds via an efficient adaptive Gibbs sampler, supporting both continuous and binary data and allowing for straightforward missing data imputation when dealing with incomplete multiway data. We discuss fundamental properties of the proposed modeling framework, providing theoretical justification. Simulation studies and applications to chemometrics and complex ecological data offer compelling evidence of its advantages over existing tensor factorization methods.
@@ 1 Introduction
In many applications, data naturally come in the form of multiway arrays or tensors such as recommender system (Bi et al., 2017), ecology (Frelat et al., 2017) and chemometrics (Tomasi and Bro, 2005). In a general setting, let $\mathcal{Y}=\{y_{i_{1}\cdots i_{K}},\ i_{k}=1,\dots,n_{k}\}$ denote a $K$-way tensor and we assume that $\mathcal{Y}\in\mathbb{R}^{n_{1}\times\dots\times n_{k}}$, with the integer $n_{k}$ being the dimension of $\mathcal{Y}$ along the $k$-th way. We are motivated by settings where $K$ is relatively small. For instance, in ecological studies examining species co-occurrence, $K=3$ with data organized in a three-way array structured as site $\times$ species $\times$ year. One naive approach to analyze this type of data is to flatten the data into a matrix and then use existing methods for matrix factorization. However, this transformation fails to preserve the data structure, making learning low-dimensional relationships more difficult.
@@
Tensor decomposition methods based on the low-rank tensor approximation (Kolda and Bader, 2009) offer a powerful tool to analyze multiway data, by providing a concise representation that effectively captures the dominant underlying features. The two main forms of tensor decomposition, which extend the matrix singular value decomposition, are the Tucker decomposition and its restriction known as the parallel factor (PARAFAC) decomposition. PARAFAC factorization decomposes a tensor into a sum of $R$ rank-one tensor, where $R$ is referred to as the rank of the tensor, while Tucker factorization decomposes a tensor into a small core tensor multiplied by a matrix along each mode. Tucker representation is more flexible than the PARAFAC, as it includes a multi-rank $(R_{1},\dots,R_{k})$ that allows for different decomposition ranks for various tensor dimension margins. This flexibility enables a more effective data compression, as not all the dimensions of a tensor might require all $R$ principal components in order to be faithfully represented by such a decomposition. In this case, PARAFAC rank can be considerably higher than the $K$ ranks composing the Tucker multi-rank. Tucker representation is particularly effective when the tensor dimensions $n_{1},\dots,n_{K}$ differ from each other.
@@
Bayesian tensor decomposition models (Chu and Ghahramani, 2009) are naturally appealing, as they allow for straightforward quantification of uncertainty in estimating model parameters, while providing a principled mechanism for dealing with missing data. Partially observed multiway data, indeed, are abundant in numerous fields; see for example Tomasi and Bro (2005) for applications in chemometrics. Missing data can arise in a variety of settings due to loss of information, errors in the data collection process, or costly experiments. The problem of recovering latent structures from incomplete tensors to reconstruct the missing entries is often referred to as tensor completion. For tensor factorization models in a Bayesian framework see for instance Hoff (2011) and Xu et al. (2013). Refer to Shi and Shen (2024) for a general review of Bayesian methods in tensor analysis, including topics beyond tensor completion, such as tensor regression models (Guhaniyogi et al., 2017; Papadogeorgou et al., 2021).
@@
An open problem in low-rank tensor factorization literature is the choice of the latent rank. Current practice typically specifies the ranks in advance or uses heuristics for choosing the ranks based on repeated fitting under different choices. However, in the Tucker case, an exhaustive search over the space of possible multi-ranks may be infeasible, since the number of possibilities increases exponentially with the order of the tensor. In principle, one can apply a matrix rank determination method for matrix separately to each mode-$k$ matricization of a tensor, such as the criteria proposed in Ahn and Horenstein (2013). However, these methods are inadequate for binary tensor data or tensors with missing values. A more elegant approach relies on Bayesian nonparametric methods with Rai et al. (2014) and Zhao et al. (2015) being interesting contributions in the PARAFAC context.
@@
Motivated by the above considerations and the need for an intermediate formulation that balances the simplicity of the PARAFAC with the flexibility of the Tucker in a parsimonious way, we introduce a flexible Bayesian Tucker decomposition model where the multi-rank is not pre-specified but is adaptively learned from the data. We rely on successful Bayesian nonparametric models that include more than enough latent elements, with their selection induced by appropriate shrinkage priors that adaptively eliminate unnecessary components. Exploiting the literature on infinite factor model (Bhattacharya and Dunson, 2011; Legramanti et al., 2020; Schiavon et al., 2022) we introduce a multiway extension of the cumulative shrinkage process (CUSP) of Legramanti et al. (2020) to apply joint shrinkage for multiple latent quantities along different dimensions. While the proposed multiway CUSP builds on the foundation of Legramanti et al. (2020), it addresses unique challenges inherent to tensor-structured data. For example, besides providing shrinkage towards low-rank decomposition of the tensor margins, the proposed method also enables local shrinkage in the latent factors by assuming a suitable shrinkage prior for the elements of the core tensor. Incorporating zero entries in the core tensor is crucial for reducing complexity, as inference in the Tucker model scales with the size of the core tensor. By shrinking towards zero the elements of the core tensor we are moving it to something similar to a superdiagonal tensor with different dimension margins, and thus bridging PARAFAC and Tucker decompositions. For a contribution in this direction for probability tensors for multivariate categorical data see the collapsed Tucker decomposition in Johndrow et al. (2017).
@@
In the next sections, we first establish the notation and then present our proposed Bayesian adaptive Tucker (AdaTuck) decomposition. This novel Bayesian approach estimates the latent multi-rank of tensors using an increasing shrinkage prior while enforcing local sparsity. We will discuss appealing theoretical properties and outline straightforward implementations for modeling both continuous and binary incomplete multiway data using fully conjugate Bayesian analysis. Posterior distributions are approximated through closed-form Gibbs sampling updates, making inference highly scalable and straightforward to predict missing entries in the data, as discussed in Section 3. The empirical performances of the proposed approach are tested against alternatives in Section 4, including simulated and real data examples from a chemometrics application and a complex ecological study. Section 5 presents the main findings of the paper and outlines possible directions for future research. Technical details, including proofs and additional numerical results are reported in the Supplementary Materials.
@@ 2.2 Bayesian model specification
Proposition 1 provides a simple and intuitive formulation for the expected value of the latent multi-rank, yielding useful insights for the choice of the hyperparameters $\alpha^{(k)}$. Thus, $\alpha^{(k)}$ could be set to the expected numbers of active components of the $k$-th mode for $k=1,\dots,K$.
@@
Let $M_{\textbf{i}}^{R_{1},\dots,R_{K}}=\sum_{\mathbf{r},r_{k}=R_{k}+1}^{\infty}g_{\mathbf{r}}\prod_{k=1}^{K}u^{(k)}_{i_{k}r_{k}}$ be the residuals for a generic index vector $\mathbf{i}$. Assume $E(g_{\mathbf{r}}^{2})\leq\zeta$ with $\zeta$ finite. Then, $\forall\epsilon>0$, $\mathrm{pr}\{(M_{\textbf{i}}^{R_{1},\dots,R_{K}})^{2}>\epsilon\}<(\zeta/\epsilon)\prod_{k=1}^{K}\alpha^{(k)}\{\alpha^{(k)}(1+\alpha^{(k)})^{-1}\}^{R_{k}}$ for each $i\in I$.
@@
The hierarchical Bayesian model specification is completed specifying a prior distribution for the error variance in (3). For simplicity we let $\sigma^{2}\sim\text{InvGa}(a_{\sigma},b_{\sigma})$. When $\mathcal{Y}$ has $m<n$ missing entries, the likelihood can be easily adapted from (3) by removing, from the covariance matrix, the associated $m$ rows and columns. In this case, our Bayesian approach seamlessly allows to perform missing data imputation by simulating from the posterior predictive distribution. Extending the cumulative shrinkage process framework to predict missing values is a key contribution that sets apart our work from the original CUSP (Legramanti et al., 2020) for matrix data. See Sections 4.1 for compelling results on tensor completion using synthetic data, as well as an image reconstruction application.
@@ 2.4 Consistency in rank selection
One of the key aspects of AdaTuck is its ability to learn the rank along each dimension, through the multiway increasing shrinkage prior. Consistent with this characteristic, in this section, we investigate the asymptotic properties of the adopted multiway CUSP in a simplified setting. This analysis extends the discussion in Section 3.3 of Kowal and Canale (2023) and relates to the problem of estimating a high-dimensional mean vector from a single multivariate observation (Castillo and van der Vaart, 2012; Ročková, 2018). While these theoretical results are derived under simplified assumptions and do not directly apply to the situations discussed in previous sections, they provide valuable insights for the proposed approach. The asymptotic analysis serves as a principled justification for the multiway CUSP’s ability to discard superfluous dimensions in controlled settings, offering theoretical support for its rank-learning capability. The empirical results presented in Section 4, which focus on the actual model discussed earlier, demonstrate that the method maintains its consistency in rank estimation, even when applied beyond the simplified framework of this section.
@@
with $\theta^{(k)}_{i}$ following the prior structure in (4). While this setting differs from the Tucker decomposition model, we observe that the above likelihood shares some similarities with the proposed models as it corresponds to the multiplicative term in the definition of the elementwise Tucker decomposition in (2). Hence $y_{i}\sim N(\eta_{i},1)$ for $i=1,2,\dots$ with $\eta_{i}=\prod_{k=1}^{K}u_{i}^{(k)}$. We assume that the true data generating process is
@@
According to Lemma 1, the prior probability that $M_{n}$ exceeds a small threshold is exponentially small. To connect the prior behavior in Lemma 1 to the posterior we denote with $\mathrm{pr}_{0}(\cdot)$ and $E_{0}(\cdot)$ the probability and the expectation respectively under the true distribution of the data, described in (10). Theorem 2 shows that the posterior probability that $M_{n}$ exceeds a shrinking $\epsilon_{n}$ for $n\to\infty$ goes to zero in expectation thus providing theoretical justification for the use of the proposed multiway CUSP to learn the ranks of model (1). Intuitively, this implies that the posterior distribution asymptotically concentrates on models where redundant parameters—those not contributing to the true data-generating structure—are discarded, ensuring consistent recovery of the tensor ranks as the sample size increases.
@@
Let $\epsilon_{n}\to 0$ with $\epsilon_{n}^{1/\tilde{R}^{0n}}>\alpha^{(k)}/(1+\alpha^{(k)})$ for any $k=1\dots,K$. We assume $A\tilde{R}^{0n}/n<1/2$ with $A>1/2$. Let $\theta_{\infty}=\theta_{\infty}(n)$ in the multiway CUSP prior (4) with $\theta_{\infty}(n)<({\tilde{R}}^{n0}/n)^{1/(2K)}$. Then, $\lim_{n\to\infty}E_{0}\{\mathrm{pr}(M_{n}>\epsilon_{n}\mid y_{1},\dots,y_{n})\}=0.$
@@ 4.1 Synthetic data
    In this section, we evaluate the performance of AdaTuck with synthetic data. Specifically, we assess its performance in terms of learning the true latent multi-rank of the Tucker factorization, under correct specification. Additionally, we assess its goodness-of-fit by randomly removing some entries of the tensor data before fitting and then imputing these missing values by exploiting the posterior predictive distribution. See Section 3 of the Supplementary Materials for additional experiments in data completion with real image data.
@@
    To assess the performance of our proposed approach in learning the true underlying multi-rank $(R_{1}^{0},R_{2}^{0},R_{3}^{0})$, we compute the Monte Carlo average over the $B$ replicates of the posterior median of $R_{1}^{*}$, $R_{2}^{*}$ and $R_{3}^{*}$ under the AdaTuck model. The results for the $B$ different datasets across the two different scenarios are shown in Figure 1. In both scenarios, our proposed method correctly recovers the true multi-rank for all tensor margins, thus efficiently learning the truncation level in the adaptive Gibbs sampler. AdaTuck posterior estimations of the multi-rank highly concentrate around the true values $(R_{1}^{0},R_{2}^{0},R_{3}^{0})$, and the results are robust with respect to the number of missing values in the data. Figure S2 in the Supplementary Materials illustrates the results for the probit-AdaTuck model in the same scenarios, showing its effectiveness in recovering the true multi-rank also in the binary data setting. For continuous tensor data, we benchmark multi-rank estimations using the Eigenvalue Ratio (ER) and the Growth Ratio (GR) criteria (Ahn and Horenstein, 2013), applied to each mode-$k$ matricization $Y(k)$. These methods perform well on fully observed tensors but fail for partially observed data, as they require complete matrices for eigendecomposition. When missing values are naively imputed using the mean, ER and GR severely underestimate the true multi-rank, as shown in Table S2 of the Supplementary Materials.
@@
    Subsequently, we evaluate the performance of the proposed method in predicting the missing values and reconstructing the original tensor for the two scenarios. We benchmark the results with the Bayesian tensor factorization method proposed in Khan and Ammad-ud din (2016) and implemented in the R package tensorBF, which targets multiway data with incomplete records. This method implements a Bayesian PARAFAC decomposition and achieves rank optimization through a sparsity parameter to remove redundant factors. We run tensorBF with the default options for the same number of iterations we used for AdaTuck. We compare the two methods by computing the Monte Carlo mean square error (MSE) in predicting the missing values with the reconstructed tensor based on the latent structure. Table 1 summarizes, for each scenario and model, the median and the interquartile range of the MSE obtained from the $B$ measures generated from different simulations, together with the running times. Such quantities rely on an R implementation run on a Apple M1 CPU laptop with 8 GB of RAM. Across all scenarios AdaTuck outperforms the competitor in predicting the missing entries of the tensor. As expected, tensorBF, which implements a PARAFAC decomposition, performs better in the scenario similar to the PARAFAC structure. Instead, AdaTuck with its greater flexibility provides good results in both scenarios. Additionally, AdaTuck is robust to variations in the amount of available data, maintaining stable performance even as the percentage of missing values increases, whereas the competitor’s performance gets worse as missing entries increase.
    Table 1: Performance of AdaTuck and TensorBf in predicting the missing entries for the 20 simulated datasets for the two scenarios. The shorthand ‘NA(%)’ represents the percentage of missing entries with respect to the total tensor entries. The runtime is in seconds.
@@
    $(R_{1}^{0},R_{2}^{0},R_{3}^{0})$ NA (%) MSE MSE runtime MSE MSE runtime
    (median) (IQR) (median) (median) (IQR) (median)
    (5,5,5) 30 0.22 0.01 1053 33.00 47.27 240
    50 0.23 0.02 962 46.73 374.78 241
    (10,7,3) 30 0.23 0.01 2234 138.27 121.13 326
    50 0.24 0.01 1781 146.28 153.58 324